OpenStudy (anonymous):

use implicit differentiation to differentiate: 1+ye^(y)=x+e^(x)

OpenStudy (anonymous):

Someone PLEASE help! I'm soo stuck right now!!!

OpenStudy (anonymous):

Oh thank you!!!

OpenStudy (whpalmer4):

March through the equation taking the derivative with respect to x of each term.

OpenStudy (anonymous):

I understand that, but what's the derivative of ye^(y)? I get taht you need to use the product rule, but I get stuck at dy/dx(e^y)!

OpenStudy (whpalmer4):

The right hand side is easy — what did you get there?

OpenStudy (anonymous):

I got 1+e^x. THat wasn't a problem! It's the other side that's got me confused!

OpenStudy (whpalmer4):

I just want to make sure that you've gotten it so I don't get an abuse claim about showing you the answer without giving you a chance to work it out.

OpenStudy (whpalmer4):

So, that gives us \[\frac{d}{dx}[1+ye^y] = 1+e^x\]What is the derivative of 1?

OpenStudy (anonymous):

0. but how do you get e^x?

OpenStudy (whpalmer4):

Excellent. So that leaves \[\frac{d}{dx}[ye^y] = 1+e^x\]That's a product there on the left side, so we can use the product rule. Do you remember what that is?

OpenStudy (anonymous):

y*(dy/dx(e^y))+e^y(dy/dx(y)) Right?

OpenStudy (whpalmer4):

Let's let \[u=e^y\] and \[v=y\] then \[\frac{d}{dx}[uv]=v\frac{du}{dx}+u\frac{dv}{dx}\]

OpenStudy (anonymous):

So I was right, right? y*(dy/dx(e^y)) + e^(y)*(dy/dx(y))

OpenStudy (whpalmer4):

I'm not 100% convinced by the way you are writing it...

OpenStudy (anonymous):

What do you mean? y multiplied by dy/dx(e^y) + e^(y) mulitplied by dy/dx(y)

OpenStudy (whpalmer4):

do you mean \[y(\frac{dy}{dx}(e^y)) + e^y(\frac{dy}{dx}(y))\]

OpenStudy (anonymous):

Exactly!! So how would I take the derivative of e^y?

OpenStudy (whpalmer4):

use the chain rule...

OpenStudy (whpalmer4):

the reason I don't like the way you wrote it is that it is unclear that you don't mean x(y) in some of those places, for example, and often in doing implicit differentiation you'll want to do things like y(x) to remind yourself that y is a function of x...

OpenStudy (anonymous):

How would I use the chain rule on e^y?

OpenStudy (whpalmer4):

Let \(u=y\) and that gives you \[\frac{d}{dx}(e^y) = \frac{de^u}{du}\frac{du}{dx}\]

OpenStudy (anonymous):

Sooo... e^y multiplied by dy/dx

OpenStudy (whpalmer4):

you're going to end up with \[\frac{dy}{dx}\]as a factor in both terms on the left hand side.

OpenStudy (anonymous):

Was I right?

OpenStudy (whpalmer4):

What is your final answer?

OpenStudy (anonymous):

y*(e^y*(dy/dx)) + (e^y)*(dy/dx)

OpenStudy (anonymous):

Just for the left hand side!! The other side we already know!

OpenStudy (anonymous):

Hey are you there? Help, please?

OpenStudy (abb0t):

wut?

OpenStudy (whpalmer4):

yeah, that looks right.

OpenStudy (anonymous):

Okayy, so then I would isolate the dy/dx! But how?

OpenStudy (whpalmer4):

\[ye^y\frac{dy}{dx}+e^y\frac{dy}{dx} = 1 + e^x\]You can factor the left side, can't you?

OpenStudy (anonymous):

Take out e^y and dy/dx, right?

OpenStudy (whpalmer4):

You could do that. At least the dy/dx, in any case.

OpenStudy (anonymous):

So the final anwer should look like: dy/dx=1+e^x/(ye^y +e^y)

OpenStudy (whpalmer4):

That should give you \[\frac{dy}{dx} (ye^y+e^y) = 1+e^x\] or \[\frac{dy}{dx} = \frac{1+e^x}{e^y(y+1)}\] or if you factored it more completely: \[e^y\frac{dy}{dx}(1+y) = 1+e^x\]\[e^y\frac{dy}{dx}=\frac{1+e^x}{1+y} \]\[\frac{dy}{dx}=\frac{e^{-y}}{1+y}+\frac{e^xe^{-y}}{1+y} = \frac{e^{-y}}{1+y} + \frac{e^{x-y}}{1+y}\]

OpenStudy (whpalmer4):

Be careful about writing divisions: you wrote dy/dx = 1+e^x/(ye^y+e^y) which equals \[\frac{dy}{dx}=1+\frac{e^x}{ye^y+e^y}\] by the rules of operator precedence...

OpenStudy (anonymous):

Perfect!! I understand Now! Thanks for all the help! You really cleared that up for me!! And the answer key they gave us has it written the way I gave it... Thanks again!! *gives cookie* :)

OpenStudy (whpalmer4):

Well, the answer key is wrong if it wrote it that way, as the two are not equivalent. You need parentheses around the numerator: dy/dx = (1+e^x)/(ye^y+e^y) At \(x =2, y = 3\), \[\frac{1+e^x}{ye^y+e^y} = \frac{1+e^2}{4e^3} \approx 0.104417\] but written carelessly, \[1+\frac{e^x}{ye^y+e^y} = 1+\frac{1}{4e} \approx 1.90197\]